OpenStudy (anonymous):

Need quick help please! Ben chooses a number from among the fist 20 counting numbers 1,2,......19,20. What is the probability that the number chosen is divisible by 4 or 5?

OpenStudy (anonymous):

count the numbers that is divisible by 4: 4, 8, 12, 16, 20 count the numbers that is divisible by 5: 5, 10, 15 add them up and divide by 20.

OpenStudy (anonymous):

it is not replaced @ParthKohli

OpenStudy (anonymous):

\[\frac{5}{20} + \frac{3}{19}= ?\]

OpenStudy (anonymous):

@dpaInc : So would it be 8/20?

OpenStudy (anonymous):

8/20=2/5... one of those...:)

OpenStudy (anonymous):

Thank you! :D